Charlie, who is just 16 years old, won a lottery. However, the lottery does not pay out money until a person is at least 25 years old. Charlie hires a financial consultant and this consultant suggests that Charlie invest the money and then begin receiving monthly payments of 2500$ once he reaches age 25. The consultant also suggests that Charlie should protect himself from inflation by having his payments increase by 2.3% each year. Thus, the 12 monthly payments in year 1 will be 2500$ (age 25 until age 25+11months), the 12 monthly payments for year 2 will be 2500(1.023)$ (age 26 to age 26+11months), the 12 payments for year 3 will be 2500(1.023)2$ and so on.(NOTE: this is a situation where payments are monthly, but the percentage increase is annually.) It turns out that Charlie has won enough money for this payment scheme to last for a total of 25 years after he turns 25. If Charlie can earn j1 = 8% on his lottery winnings and the consultant charges 1.5% of the winnings as a fee (payable upfront when charlie is 16), how much did Charlie win in the lottery?
Answer #1

the rate of interest given is 8% as per present value formula Present value = Amount/(1+r)^n, where n = no. of years and r = rate

so, The discounting factor (for 1 $) would be = 1/(1+8%)^n ,

Note : In the question, no mention about the fraction values to be used for decimal places. so for the convenience of the calculation, the intermediary calculations are all rounded up to nearest $.

and also assumed that charlie just got attained the age of 16 years.

Monthly Payment 26316 calculation of Present value of payments. for 25 years Monthly Total Monthly Discourting Discosted Year

Let us assume the winning amount of charlie be d e So, the upfront fel = 1.5% of = 0.015a . So, what he would have got = x -0

So, the amount required to Pay upfront fee we lottery - 195400 amount remaining l.e. a-0.0158 - 195400 0.985x 195400 se 19540